OpenStudy (anonymous):

What is the value of g[f(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

you can start by evaluating the inside part, f(4). If f(x) = 2x+1, then what is f(4)?

OpenStudy (anonymous):

is the answer 3?

OpenStudy (anonymous):

When I find f(4), I get f(4) = 2(4)+1 = 9 So g[ f(4) ] = g( 9 ) since f(4) is the same as 9. And you know g(x) = 2x - 5 So g(9) = ?